OpenStudy (anonymous):

For the function f(x)=4/x+3. Find the average rate of change of f from 1 to x.

OpenStudy (anonymous):

really from 1 to \(x\) ?

OpenStudy (anonymous):

i guess it is \[\frac{1}{x-1}\int_1^x\frac{4}{t+3}dt\]

OpenStudy (anonymous):

anti derivative is \(4\ln(t+3)\) so integral would be \[\frac{1}{x-1}4\left(\ln(x+3)-\ln(4)\right)\]

OpenStudy (anonymous):

So how would I find the rate of change?

OpenStudy (anonymous):

what class is this for? it it calculus?

OpenStudy (anonymous):

College Algebra

OpenStudy (anonymous):

then my answer is not the one you are looking for

OpenStudy (anonymous):

are you sure it is the rate of change from 1 to \(x\) and not from 1 to some number

OpenStudy (anonymous):

Blahhh so hard for me unfortunately

OpenStudy (anonymous):

I am sure

OpenStudy (anonymous):

okay then this is what you need to compute we can do it, it is not that hard \[\frac{f(x)-f(1)}{x-1}\]

OpenStudy (anonymous):

this is the change in \(f\) over the change in \(x\) just like the slope of a line now comes some annoying algebra

OpenStudy (anonymous):

They gave me options for the answer but I dont know how to put fractions on this site

OpenStudy (anonymous):

no problem, we can work out the answer first off we need \(f(1)=\frac{4}{1+3}=\frac{4}{4}=1\)

OpenStudy (anonymous):

so now we have to compute \[\frac{f(x)-1}{x-1}\]

OpenStudy (anonymous):

we get \[\frac{\frac{4}{x+3}-1}{x-1}\] as a first step

OpenStudy (anonymous):

a) -1/x+3 b) 1/x+3 c) 4/x(x+3) d) 4/(x-1)(x+3)

OpenStudy (anonymous):

now multiply top and bottom by \(x+3\) to clear the fraction you get \[\frac{\frac{4}{x+3}-1}{x-1}\times \frac{x+3}{x+3}\] \[=\frac{4-(x+3)}{(x-1)(x+3)}\] \[=\frac{4-x-3}{(x-1)(x+3)}\] \[=\frac{1-x}{(x+3)}\]

OpenStudy (anonymous):

typo on the last line

OpenStudy (anonymous):

it should be \[=\frac{1-x}{(x-1)(x+3)}\]

OpenStudy (anonymous):

then since \(\frac{1-x}{x-1}=-1\) the final answer is \[\frac{-1}{x+3}\]

OpenStudy (anonymous):

Thank you!!! ;)